### v1

# Joint Time–Position Statistics and Fisher Information in Drift–Diffusion Molecular Channels

Abstract
This letter presents a closed-form characterization of the joint distribution of first arrival time (FAT) and first arrival position (FAP) in diffusion-based molecular communication (MC) systems with drift. Prior studies have investigated FAT modeling via inverse Gaussian distributions [1] and applied FAT statistics for parameter estimation and synchronization tasks [2, 3], while more recent work has characterized FAP for spatial channel analysis [4]. In contrast, we derive an explicit joint probability density function (PDF) under constant drift and isotropic diffusion in arbitrary spatial dimensions. Our result reveals a nontrivial coupling between arrival time and lateral position, generalizing known inverse Gaussian models. We further compute the Fisher information matrix (FIM) with respect to key channel parameters, showing that the joint observation enables estimation of lateral drift and improves sensitivity to the diffusion coefficient—capabilities not achievable with time-only or position-only models. This joint framework enhances the modeling and inference capabilities for molecular communication channels where spatial randomness itself carries non-negligible information.
Index Terms:
Molecular communication (MC), drift-diffusion channels, first arrival distribution, joint time-position distribution, Fisher information.
I Introduction
Molecular communication (MC) via diffusion is a biologically inspired paradigm for nanoscale information exchange, where molecules propagate through random motion under background drift [5, 6]. A fundamental performance metric in such systems is the first arrival time (FAT), which governs detection latency and is well modeled by the inverse Gaussian (IG) distribution [7] in one-dimensional drift–diffusion channels [1]. Beyond modeling, FAT statistics have been widely employed for synchronization and parameter estimation tasks [2, 3].
In higher-dimensional settings, however, molecules may reach different spatial locations on the absorbing receiver due to lateral Brownian motion. This motivates the study of the first arrival position (FAP), which complements the temporal information provided by FAT. Prior work [4] derived a closed-form FAP distribution using a stochastic PDE framework, demonstrating its role in channel capacity analysis and its applicability to arbitrary spatial dimensions and drift directions.
Most existing studies, however, have focused on either FAT or FAP in isolation. In contrast, this letter develops a joint characterization of FAT and FAP under constant drift and isotropic diffusion. We derive a closed-form joint PDF in $D$ dimensions, revealing explicit spatiotemporal coupling. Related joint designs have been explored in diffusion channels that combine time with other modulation dimensions [8].
We further analyze the Fisher information matrix (FIM) associated with the joint observation, quantifying its utility for parameter estimation. Prior work has applied FIM analysis to other MC settings [9, 2, 3], but not in conjunction with FAT–FAP statistics. Our results show that the joint model enables identification of lateral drift components and improves sensitivity to diffusivity—capabilities not achievable with FAT-only or FAP-only models.
This joint perspective provides a stronger foundation for inference and system design in MC applications, particularly in regimes where spatial randomness carries non-negligible information.
II System Model
We consider a single-input single-output (SISO) molecular communication (MC) system in a $D$-dimensional unbounded fluid medium with constant drift and diffusion. The transmitter (TX) and receiver (RX) are modeled as parallel (D1)(D-1)-dimensional hyperplanes, separated by one unit length along the $x^{(1)}$ axis.
To simplify the analysis, we normalize length by the TX–RX distance $\lambda$ and time by $\lambda/v_{p}$, where $v_{p}>0$ denotes the perpendicular drift component. Under this normalization, the system becomes dimensionless; in particular, the perpendicular drift is scaled to $1$, so that the drift vector takes the form $\mathbf{v}=(1,v^{(2)},\ldots,v^{(D)})$.
The TX emits a single molecule from location
$\mathbf{x}_{0}=(0,x_{0}^{(2)},\ldots,x_{0}^{(D)})$ (1)
at time $t=0$, and the RX fully absorbs the molecule upon its first arrival at the hyperplane $\{x^{(1)}=1\}$.
The molecule propagates under constant drift and isotropic diffusion, and its position $\mathbf{X}_{t}=(X_{t}^{(1)},\ldots,X_{t}^{(D)})$ evolves according to the stochastic differential equation
$\mathrm{d}\mathbf{X}_{t}=\mathbf{v}\,\mathrm{d}t+\sigma\,\mathrm{d}\mathbf{B}_{t},$ (2)
where $\mathbf{v}=(1,v^{(2)},\ldots,v^{(D)})\in\mathbb{R}^{D}$ is the drift vector, $\sigma>0$ is the dimensionless diffusion coefficient, and $\{\mathbf{B}_{t}\}_{t\geq 0}$ is a standard $D$-dimensional Wiener process with independent components.
We define the first arrival time as
$T:=\inf\{t>0:X_{t}^{(1)}=1\},$ (3)
and the first arrival position as the random vector
$\mathbf{X}_{T}=(1,X_{T}^{(2)},\ldots,X_{T}^{(D)}).$ (4)
In the next section, we derive the joint distribution of $(T,\mathbf{X}_{T})$, which captures the spatiotemporal coupling of first-arrival dynamics under drift–diffusion.
III Joint Time-Position Distribution of First Arrivals
We derive the joint probability density function (PDF) of the first arrival time $T$ and the first arrival position $\mathbf{X}_{T}$ of a diffusing particle under constant drift in a dimensionless setting. The transmitter is located on the hyperplane $\{x^{(1)}=0\}$, and the fully absorbing receiver lies on $\{x^{(1)}=1\}$. A molecule is released from
$\mathbf{x}_{0}=(0,x_{0}^{(2)},\ldots,x_{0}^{(D)})$ (5)
at time $t=0$, and its position $\mathbf{X}_{t}$ evolves according to the stochastic differential equation
$\mathrm{d}\mathbf{X}_{t}=\mathbf{v}\,\mathrm{d}t+\sigma\,\mathrm{d}\mathbf{B}_{t},$ (6)
where $\mathbf{v}=(1,v^{(2)},\ldots,v^{(D)})\in\mathbb{R}^{D}$ is the drift vector, $\sigma>0$ is the diffusion coefficient, and $\{\mathbf{B}_{t}\}_{t\geq 0}$ is a standard $D$-dimensional Wiener process with independent components.
The first arrival time is defined as
$T:=\inf\{t>0:X_{t}^{(1)}=1\},$ (7)
and the corresponding first arrival position is $\mathbf{X}_{T}=(1,X_{T}^{(2)},\ldots,X_{T}^{(D)})$. Throughout, we assume that the components of $\mathbf{B}_{t}$ are independent and that the receiver is absorbing only in the $x^{(1)}$-direction.
For notational convenience, we denote by $\mathbf{u}^{(a:b)}:=(u^{(a)},u^{(a+1)},\ldots,u^{(b)})$ the subvector of $\mathbf{u}$ from the $a$-th to the $b$-th component.
To justify the forthcoming factorization, note that for any $t>0$ and $\mathbf{x}=(1,\mathbf{x}^{(2:D)})$, the joint density of $(T,\mathbf{X}_{T})$ can be decomposed as
$\begin{split}f_{T,\mathbf{X}_{T}}(t,\mathbf{x})&\overset{(a)}{=}f_{T}(t)\,f_{\mathbf{X}_{T}|T}(\mathbf{x}|t)\\ &\overset{(b)}{=}f_{T}(t)\,f_{\mathbf{X}_{t}|T}(\mathbf{x}|t)\\ &\overset{(c)}{=}f_{T}(t)\,\delta(x^{(1)}-1)\,f_{\mathbf{X}_{t}^{(2:D)}}(\mathbf{x}^{(2:D)}),\end{split}$ (8)
where (a) follows from the definition of the conditional density, (b) uses the equality in distribution $(\mathbf{X}_{T}|T=t)\stackrel{{\scriptstyle d}}{{=}}(\mathbf{X}_{t}|T=t)$, and (c) holds because the lateral process $\mathbf{X}^{(2:D)}_{t}:=(X_{t}^{(2)},\ldots,X_{t}^{(D)})$ is independent of $T$. This independence stems from the geometry of the problem: the hitting time $T$ is determined solely by the one-dimensional process $\{X_{t}^{(1)}\}_{t\geq 0}$, while the components of the Wiener process are mutually independent. This argument extends the proof of [10, Thm. 2.37] to arbitrary dimension $D$.
Therefore, the joint density can be factorized as
$f_{T,\mathbf{X}_{T}}(t,\mathbf{x})=f_{T}(t)\,\delta(x^{(1)}-1)\,f_{\mathbf{X}_{t}^{(2:D)}}(\mathbf{x}^{(2:D)}),$ (9)
where $f_{T}(t)$ is the inverse Gaussian (IG) density
$f_{T}(t)=\frac{1}{\sqrt{2\pi\sigma^{2}t^{3}}}\exp\!\left(-\frac{(1-t)^{2}}{2\sigma^{2}t}\right)\,\mathbbm{I}\{t>0\},$ (10)
and $f_{\mathbf{X}_{t}^{(2:D)}}$ is the density of a (D1)(D-1)-dimensional Gaussian random vector:
$f_{\mathbf{X}_{t}^{(2:D)}}(\mathbf{x}^{(2:D)})=\frac{1}{(2\pi\sigma^{2}t)^{(D-1)/2}}\\[-4.30554pt] \times\exp\!\Biggl{(}-\frac{\bigl{\|}\mathbf{x}^{(2:D)}-\mathbf{x}_{0}^{(2:D)}-\mathbf{v}^{(2:D)}t\bigr{\|}^{2}}{2\sigma^{2}t}\Biggr{)}.$ (11)
Combining (10) and (11), we obtain
$\begin{split}f_{T,\mathbf{X}_{T}}(t,\mathbf{x})&=\delta(x^{(1)}-1)\,(2\pi\sigma^{2})^{-D/2}\,t^{-D/2-1}\,\mathbbm{I}\{t>0\}\\ \times\exp&\!\Biggl{(}-\frac{(1-t)^{2}+\bigl{\|}\mathbf{x}^{(2:D)}-\mathbf{x}_{0}^{(2:D)}-\mathbf{v}^{(2:D)}t\bigr{\|}^{2}}{2\sigma^{2}t}\Biggr{)}.\end{split}$ (12)
The marginal density of the first arrival position (FAP) is obtained by integrating over $t$:
$f_{\mathbf{X}_{T}}(\mathbf{x})=\int_{0}^{\infty}f_{T,\mathbf{X}_{T}}(t,\mathbf{x})\,dt.$ (13)
Identifying the resulting integral as a special case of the identity involving the modified Bessel function $K_{\nu}(\cdot)$ [11, Eq. (3.478.4)], we obtain
$\begin{split}f_{\mathbf{X}_{T}}(\mathbf{x})&=\delta(x^{(1)}-1)\,2\left(\frac{\|\mathbf{v}\|}{2\pi\sigma^{2}}\right)^{D/2}\exp\!\left(\frac{\mathbf{v}\cdot(\mathbf{x}-\mathbf{x}_{0})}{\sigma^{2}}\right)\\ &\quad\times\frac{K_{D/2}\!\left(\frac{\|\mathbf{v}\|\,\|\mathbf{x}-\mathbf{x}_{0}\|}{\sigma^{2}}\right)}{\|\mathbf{x}-\mathbf{x}_{0}\|^{D/2}}.\end{split}$ (14)
This expression generalizes earlier results for $D=2,3$ [4] and provides a closed-form characterization of the FAP distribution in arbitrary spatial dimensions under drift–diffusion transport.
IV Fisher Information Analysis
We next analyze the information contained in the joint observation $(T,\mathbf{X}_{T})$ with respect to key model parameters. In particular, we compute the Fisher information matrix (FIM) associated with the parameter vector $\boldsymbol{\theta}=(\sigma,v^{(2)},\ldots,v^{(D)})$, based on the joint PDF derived in Section III. The FIM characterizes the fundamental limits of parameter estimation accuracy for drift and diffusion in molecular communication systems.
From (12), the joint density can be written as
$f_{T,\mathbf{X}_{T}}(t,\mathbf{x})=\delta(x^{(1)}-1)\,(2\pi\sigma^{2})^{-D/2}\,t^{-D/2-1}\,\mathbbm{I}\{t>0\}$
$\quad\quad\times\exp\!\left(-\frac{(1-t)^{2}+\|\mathbf{x}^{(2:D)}-\mathbf{x}_{0}^{(2:D)}-\mathbf{v}^{(2:D)}t\|^{2}}{2\sigma^{2}t}\right).$ (15)
The corresponding log-likelihood function (up to the support constraint $t>0$ and $x^{(1)}=1$) is
$\begin{split}&\log f_{T,\mathbf{X}_{T}}(t,\mathbf{x};\boldsymbol{\theta})=-\tfrac{D}{2}\,\log(2\pi\sigma^{2})-\left(\tfrac{D}{2}+1\right)\log t\\ &\quad\quad\quad\quad-\frac{(1-t)^{2}+\|\mathbf{x}^{(2:D)}-\mathbf{x}_{0}^{(2:D)}-\mathbf{v}^{(2:D)}t\|^{2}}{2\sigma^{2}t}.\end{split}$ (16)
For convenience, define the residual vector
$\boldsymbol{\Delta}_{t}:=\mathbf{x}^{(2:D)}-\mathbf{x}_{0}^{(2:D)}-\mathbf{v}^{(2:D)}t,$ (17)
and denote its $k$-th component by $\Delta_{t}^{(k)}$ for $k\in\{2,\ldots,D\}$.
IV-A Partial Derivatives
The partial derivatives of the log-likelihood with respect to the parameters are as follows.
Derivative with respect to $\sigma$
$\frac{\partial}{\partial\sigma}\log f=-\frac{D}{\sigma}+\frac{1}{\sigma^{3}t}\Bigl{[}(1-t)^{2}+\|\boldsymbol{\Delta}_{t}\|^{2}\Bigr{]}.$ (18)
Derivative with respect to $v^{(k)}$, $k\in\{2,\ldots,D\}$
$\frac{\partial}{\partial v^{(k)}}\log f=\frac{\Delta_{t}^{(k)}}{\sigma^{2}}.$ (19)
These expressions will be used in the next subsection to compute the Fisher information matrix entries.
IV-B Fisher Information Matrix Entries
For the parameter vector $\boldsymbol{\theta}=(\sigma,v^{(2)},\ldots,v^{(D)})$, the Fisher information matrix (FIM) $\mathbf{I}(\boldsymbol{\theta})$ is defined entrywise as
$[\mathbf{I}(\boldsymbol{\theta})]_{i,j}=\mathbb{E}_{(T,\mathbf{X}_{T})}\!\left[\frac{\partial\log f}{\partial\theta_{i}}\frac{\partial\log f}{\partial\theta_{j}}\right],$ (20)
where the expectation is taken with respect to the joint distribution in (12). Since the log-likelihood is smooth in $(t,\mathbf{x},\boldsymbol{\theta})\in(0,\infty)\times\mathbb{R}^{D-1}\times(0,\infty)\times\mathbb{R}^{D-1}$, each entry also admits the equivalent form
$[\mathbf{I}(\boldsymbol{\theta})]_{i,j}=-\mathbb{E}_{(T,\mathbf{X}_{T})}\!\left[\frac{\partial^{2}\log f}{\partial\theta_{i}\partial\theta_{j}}\right].$ (21)
We first consider the entry corresponding to the diffusion coefficient $\sigma$. Differentiating the log-likelihood gives
$\frac{\partial^{2}}{\partial\sigma^{2}}\log f=\frac{D}{\sigma^{2}}-\frac{3}{\sigma^{4}t}\Big{(}(1-t)^{2}+\|\boldsymbol{\Delta}_{t}\|^{2}\Big{)}.$ (22)
Taking expectations yields
$I_{\sigma\sigma}$ $=-\mathbb{E}\!\left[\frac{D}{\sigma^{2}}-\frac{3}{\sigma^{4}T}\Big{(}(1-T)^{2}+\|\boldsymbol{\Delta}_{T}\|^{2}\Big{)}\right]$
$=\frac{3}{\sigma^{4}}\,\mathbb{E}\!\left[\frac{(1-T)^{2}}{T}+\frac{\|\boldsymbol{\Delta}_{T}\|^{2}}{T}\right]-\frac{D}{\sigma^{2}}.$ (23)
The two expectations can be computed separately. Since $T\sim\mathrm{IG}(1,1/\sigma^{2})$ with moments $\mathbb{E}[T]=1$ and $\mathbb{E}[T^{-1}]=1+\sigma^{2}$, it follows that
$\mathbb{E}_{T}\!\left[\frac{(1-T)^{2}}{T}\right]=\sigma^{2}.$ (24)
Moreover, conditioned on $T=t$, the lateral displacement satisfies $\boldsymbol{\Delta}_{T}\sim\mathcal{N}(0,\sigma^{2}tI_{D-1})$, which gives
$\mathbb{E}_{\mathbf{X}_{T}|T}\!\left[\|\boldsymbol{\Delta}_{T}\|^{2}\,\big{|}\,T\right]=(D-1)\sigma^{2}T.$ (25)
Taking expectation over $T$ yields
$\mathbb{E}_{T}\!\left[\frac{1}{T}\,\mathbb{E}_{\mathbf{X}_{T}|T}[\|\boldsymbol{\Delta}_{T}\|^{2}|T]\right]=(D-1)\sigma^{2},$ (26)
and combining the results gives
$I_{\sigma\sigma}=\frac{2D}{\sigma^{2}}.$ (27)
Turning to the cross terms between $\sigma$ and a lateral drift $v^{(k)}$, $k\in\{2,\ldots,D\}$, the mixed derivative is
$\frac{\partial^{2}}{\partial\sigma\partial v^{(k)}}\log f=-\frac{2}{\sigma^{3}}\,\Delta_{t}^{(k)}.$ (28)
Taking expectations, we obtain
$I_{\sigma,v^{(k)}}=\frac{2}{\sigma^{3}}\,\mathbb{E}[\Delta_{T}^{(k)}]=0,$ (29)
showing that the FIM remains diagonal.
Finally, for each drift component $v^{(k)}$, the second derivative takes the form
$\frac{\partial^{2}}{\partial[v^{(k)}]^{2}}\log f=-\frac{t}{\sigma^{2}},$ (30)
which implies
$I_{v^{(k)}v^{(k)}}=\frac{1}{\sigma^{2}}\,\mathbb{E}[T]=\frac{1}{\sigma^{2}}.$ (31)
In summary, the Fisher information matrix is diagonal, with diffusion contributing $I_{\sigma\sigma}=2D/\sigma^{2}$ and each lateral drift parameter contributing $I_{v^{(k)}v^{(k)}}=1/\sigma^{2}$. This shows that all drift components contribute symmetrically, while the diffusion parameter aggregates information from both temporal and spatial observations.
Figure 1: Empirical conditional joint PDF of $(T,X_{T}^{(2)})$ obtained from particle-based simulations.
Figure 2: Analytical conditional joint PDF of $(T,X_{T}^{(2)})$ computed from the closed-form expression.
IV-C Discussion
The closed-form expressions above show that the joint observation of first arrival time $T$ and position $\mathbf{X}_{T}$ provides valuable information on both drift and diffusion. Under pure vertical drift, the Fisher information $I_{v^{(k)}v^{(k)}}=1/\sigma^{2}$ is strictly positive, implying that lateral drift components $v^{(k)}$ are estimable from $\mathbf{X}_{T}$ but not from $T$ alone. Moreover, the diffusion term $I_{\sigma\sigma}$ reflects contributions from both temporal and spatial observations, indicating that the joint model carries richer information than either marginal model.
Thus, the joint spatiotemporal framework captures a more complete statistical structure than time-only or position-only observations. In practice, combining $T$ and $\mathbf{X}_{T}$ in estimation or detection can improve sensitivity to environmental parameters such as background flow direction or diffusivity heterogeneity.
V Numerical Simulations
We perform numerical experiments to validate the theoretical channel model in (12). For illustration, we focus on the two-dimensional (2D) case with a linear absorbing receiver modeled as the infinite line $\{(x^{(1)},x^{(2)})\in\mathbb{R}^{2}:x^{(1)}=1\}$. In this setting, the arrival position always satisfies $X_{T}^{(1)}=1$, so it suffices to examine the joint distribution of $(T,X_{T}^{(2)})$.
[…]
Since the simulation time horizon is limited to $2$ seconds, the empirical joint PDF obtained should be interpreted as conditional, i.e., conditioned on arrival within the interval $[0,2]$ seconds. For a fair comparison, we therefore also compute the corresponding conditional version of the theoretical joint PDF from (12). The results, shown in Fig. 2 and Fig. 2, demonstrate excellent agreement between theory and simulation.
VI Conclusion
We presented a closed-form expression for the joint distribution of the first arrival time and position of a molecule propagating in a drift–diffusion channel. This result generalizes the classical inverse Gaussian model by explicitly capturing the spatiotemporal coupling of molecular transport in arbitrary spatial dimensions.
Building on this joint distribution, we derived the Fisher information associated with key environmental parameters. The analysis shows that combining temporal and spatial observations enables estimation of lateral drift components and improves sensitivity to the diffusion coefficient—capabilities not attainable with time-only models.
Future work may relax the constant-drift assumption to more general non-uniform drift fields, where drift strength or direction varies spatially.