### v2

# Joint Time-Position Statistics and Fisher Information in Drift-Diffusion Molecular Channels

Abstract
This letter derives a closed-form joint distribution of the first arrival time (FAT) and first arrival position (FAP) in diffusion-based molecular communication (MC) systems with drift. Unlike prior work that studied FAT or FAP separately, we obtain an explicit joint probability density function under constant drift and isotropic diffusion in arbitrary dimensions, revealing a nontrivial spatiotemporal coupling. Based on this result, we compute the Fisher information matrix (FIM) and show that joint observations enable estimation of lateral drift and improve sensitivity to diffusion—capabilities not attainable with time-only or position-only models.
Index Terms:
Molecular communication (MC), drift-diffusion channels, first arrival distribution, joint time-position distribution, Fisher information.
I Introduction
Molecular communication (MC) via diffusion is a biologically inspired paradigm for nanoscale information exchange, where molecules propagate through random motion under background drift [1, 2]. A fundamental performance metric in such systems is the first arrival time (FAT), which governs detection latency and is well modeled by the inverse Gaussian (IG) distribution [3] in one-dimensional drift–diffusion channels [4]. Beyond modeling, FAT statistics have been widely employed for synchronization and parameter estimation tasks [5, 6].
In higher-dimensional settings, however, molecules may reach different spatial locations on the absorbing receiver due to lateral Brownian motion. This motivates the study of the first arrival position (FAP), which complements the temporal information provided by FAT. Prior work [7] derived a closed-form FAP distribution using a stochastic PDE framework, demonstrating its role in channel capacity analysis and its applicability to arbitrary spatial dimensions and drift directions.
Most existing studies, however, have focused on either FAT or FAP in isolation. In contrast, this letter develops a joint characterization of FAT and FAP under constant drift and isotropic diffusion. We derive a closed-form joint PDF in $D$ dimensions, revealing explicit spatiotemporal coupling. Related joint designs have been explored in diffusion channels that combine time with other modulation dimensions [8].
We further analyze the Fisher information matrix (FIM) associated with the joint observation, quantifying its utility for parameter estimation. Prior work has applied FIM analysis to other MC settings [9, 5, 6], but not in conjunction with FAT–FAP statistics. Our results show that the joint model enables identification of lateral drift components and improves sensitivity to diffusivity estimation—capabilities not achievable with FAT-only or FAP-only models.
This joint perspective provides a stronger foundation for parameter estimation and system design in MC applications, particularly in regimes where spatial randomness carries non-negligible information.
II System Model
We consider a single-input single-output (SISO) molecular communication (MC) system in a $D$-dimensional unbounded fluid medium with constant drift and diffusion, where a transmitter (TX) releases molecules that are absorbed by a receiver (RX). We adopt a non-dimensionalized formulation: length is normalized by the transmitter–receiver (TX–RX) distance $\lambda$, and time is normalized by $\lambda/v_{p}$, where $v_{p}>0$ denotes the perpendicular [^1] Namely, $v_{p}$ is perpendicular to the parallel hyperplanes TX and RX. drift component. Under this scaling, the system becomes dimensionless; in particular, the perpendicular drift is scaled to $1$, so that the drift vector takes the form $\mathbf{v}=(1,v^{(2)},\ldots,v^{(D)})$. Unless otherwise noted, all theoretical results in this section are expressed in these normalized units.
After normalization, the TX and RX are modeled as parallel (D1)(D-1)-dimensional hyperplanes separated by one unit length along the $x^{(1)}$ axis. The TX emits a single molecule from location
$\mathbf{x}_{0}=(0,x_{0}^{(2)},\ldots,x_{0}^{(D)})$ (1)
at time $t=0$, and the RX fully absorbs the molecule upon its first arrival at the hyperplane $\{x^{(1)}=1\}$.
The molecule propagates under constant drift and isotropic diffusion, and its position $\mathbf{X}_{t}=(X_{t}^{(1)},\ldots,X_{t}^{(D)})$ evolves according to the stochastic differential equation
$\mathrm{d}\mathbf{X}_{t}=\mathbf{v}\,\mathrm{d}t+\sigma\,\mathrm{d}\mathbf{B}_{t},$ (2)
where $\sigma>0$ is the dimensionless diffusion coefficient and $\{\mathbf{B}_{t}\}_{t\geq 0}$ is a standard $D$-dimensional Wiener process with independent components.
We define the first arrival time as
$T:=\inf\{t>0:X_{t}^{(1)}=1\},$ (3)
and the corresponding first arrival position as
$\mathbf{X}_{T}=(1,X_{T}^{(2)},\ldots,X_{T}^{(D)}).$ (4)
III Joint Time-Position Distribution of First Arrivals
We now derive the joint probability density function (PDF) of the first arrival time $T$ and first arrival position $\mathbf{X}_{T}$ under the drift–diffusion dynamics defined in Section II.
For notational convenience, we denote by $\mathbf{u}^{(a:b)}:=(u^{(a)},u^{(a+1)},\ldots,u^{(b)})$ the subvector of $\mathbf{u}$ from the $a$-th to the $b$-th component.
To justify the forthcoming factorization, note that for any $t>0$ and $\mathbf{x}=(1,\mathbf{x}^{(2:D)})$, the joint density of $(T,\mathbf{X}_{T})$ can be decomposed as
$\begin{split}f_{T,\mathbf{X}_{T}}(t,\mathbf{x})&=f_{T}(t)\,f_{\mathbf{X}_{T}|T}(\mathbf{x}|t)\\ &=f_{T}(t)\,f_{\mathbf{X}_{t}|T}(\mathbf{x}|t)\\ &=f_{T}(t)\,\delta(x^{(1)}-1)\,f_{\mathbf{X}_{t}^{(2:D)}}(\mathbf{x}^{(2:D)}),\end{split}$ (5)
where the last equality holds because the lateral process $\mathbf{X}^{(2:D)}_{t}:=(X_{t}^{(2)},\ldots,X_{t}^{(D)})$ is independent of $T$. The hitting time is determined solely by $\{X_{t}^{(1)}\}$, while the Wiener components are mutually independent. This extends the argument of [10, Thm. 2.37] to arbitrary dimension $D$.
Therefore, the joint density can be factorized as
$f_{T,\mathbf{X}_{T}}(t,\mathbf{x})=f_{T}(t)\,\delta(x^{(1)}-1)\,f_{\mathbf{X}_{t}^{(2:D)}}(\mathbf{x}^{(2:D)}),$ (6)
where $f_{T}(t)$ is the $\text{IG}(1,1/\sigma^{2})$ density
$f_{T}(t)=\frac{1}{\sqrt{2\pi\sigma^{2}t^{3}}}\exp\!\left(-\frac{(1-t)^{2}}{2\sigma^{2}t}\right)\,\mathbbm{I}\{t>0\},$ (7)
and $f_{\mathbf{X}_{t}^{(2:D)}}$ is the density of a (D1)(D-1)-dimensional Gaussian random vector:
$f_{\mathbf{X}_{t}^{(2:D)}}(\mathbf{x}^{(2:D)})=\frac{1}{(2\pi\sigma^{2}t)^{(D-1)/2}}\\[-4.30554pt] \times\exp\!\Biggl{(}-\frac{\bigl{\|}\mathbf{x}^{(2:D)}-\mathbf{x}_{0}^{(2:D)}-\mathbf{v}^{(2:D)}t\bigr{\|}^{2}}{2\sigma^{2}t}\Biggr{)}.$ (8)
Combining (7) and (8), we obtain
$\begin{split}f_{T,\mathbf{X}_{T}}(t,\mathbf{x})&=\delta(x^{(1)}-1)\,(2\pi\sigma^{2})^{-D/2}\,t^{-D/2-1}\,\mathbbm{I}\{t>0\}\\ \times\exp&\!\Biggl{(}-\frac{(1-t)^{2}+\bigl{\|}\mathbf{x}^{(2:D)}-\mathbf{x}_{0}^{(2:D)}-\mathbf{v}^{(2:D)}t\bigr{\|}^{2}}{2\sigma^{2}t}\Biggr{)}.\end{split}$ (9)
The marginal density of FAP is obtained as
$f_{\mathbf{X}_{T}}(\mathbf{x})=\int_{0}^{\infty}f_{T,\mathbf{X}_{T}}(t,\mathbf{x})\,dt.$ (10)
By applying the identity for the order-$\nu$ modified Bessel function $K_{\nu}(\cdot)$ [11, Eq. (3.478.4)], we obtain
$\begin{split}f_{\mathbf{X}_{T}}(\mathbf{x})&=\delta(x^{(1)}-1)\,2\left(\frac{\|\mathbf{v}\|}{2\pi\sigma^{2}}\right)^{D/2}\exp\!\left(\frac{\mathbf{v}\cdot(\mathbf{x}-\mathbf{x}_{0})}{\sigma^{2}}\right)\\ &\quad\times\frac{K_{D/2}\!\left(\frac{\|\mathbf{v}\|\,\|\mathbf{x}-\mathbf{x}_{0}\|}{\sigma^{2}}\right)}{\|\mathbf{x}-\mathbf{x}_{0}\|^{D/2}}.\end{split}$ (11)
This expression generalizes earlier results for $D=2,3$ [7] and provides a closed-form characterization of the FAP distribution in arbitrary spatial dimensions under drift–diffusion transport.
IV Fisher Information Analysis
We next analyze the information contained in the joint observation $(T,\mathbf{X}_{T})$ with respect to key model parameters. In particular, we compute the FIM associated with the parameter vector $\boldsymbol{\theta}=(\sigma,v^{(2)},\ldots,v^{(D)})$, based on the joint PDF derived in Section III. The FIM characterizes the fundamental limits of parameter estimation accuracy for drift and diffusion in molecular communication systems.
The log-likelihood function (up to the support constraint $t>0$ and $x^{(1)}=1$) corresponding to the joint PDF (9) is
$\begin{split}&\log f_{T,\mathbf{X}_{T}}(t,\mathbf{x};\boldsymbol{\theta})=-\tfrac{D}{2}\,\log(2\pi\sigma^{2})-\left(\tfrac{D}{2}+1\right)\log t\\ &\quad\quad\quad\quad-\frac{(1-t)^{2}+\|\mathbf{x}^{(2:D)}-\mathbf{x}_{0}^{(2:D)}-\mathbf{v}^{(2:D)}t\|^{2}}{2\sigma^{2}t}.\end{split}$ (12)
For convenience, define the residual vector at time $t$ as
$\boldsymbol{\Delta}_{t}:=\mathbf{x}^{(2:D)}-\mathbf{x}_{0}^{(2:D)}-\mathbf{v}^{(2:D)}t,$ (13)
and denote its $k$-th component by $\Delta_{t}^{(k)}$ for $k\in\{2,\ldots,D\}$.
IV-A Partial Derivatives
The partial derivatives of the log-likelihood with respect to the parameters are as follows.
Derivative with respect to $\sigma$
$\frac{\partial}{\partial\sigma}\log f=-\frac{D}{\sigma}+\frac{1}{\sigma^{3}t}\Bigl{[}(1-t)^{2}+\|\boldsymbol{\Delta}_{t}\|^{2}\Bigr{]}.$ (14)
Derivative with respect to $v^{(k)}$, $k\in\{2,\ldots,D\}$
$\frac{\partial}{\partial v^{(k)}}\log f=\frac{\Delta_{t}^{(k)}}{\sigma^{2}}.$ (15)
These expressions will be used in the next subsection to compute the Fisher information matrix entries.
IV-B Fisher Information Matrix Entries
For the parameter vector $\boldsymbol{\theta}=(\sigma,v^{(2)},\ldots,v^{(D)})$, the Fisher information matrix (FIM) $\mathbf{I}(\boldsymbol{\theta})$ is defined entrywise as
$[\mathbf{I}(\boldsymbol{\theta})]_{i,j}=\mathbb{E}_{(T,\mathbf{X}_{T})}\!\left[\frac{\partial\log f}{\partial\theta_{i}}\frac{\partial\log f}{\partial\theta_{j}}\right],$ (16)
where the expectation is taken with respect to the joint distribution in (9). Since the log-likelihood is smooth in $(t,\mathbf{x},\boldsymbol{\theta})\in(0,\infty)\times\mathbb{R}^{D-1}\times(0,\infty)\times\mathbb{R}^{D-1}$, each entry also admits the equivalent form
$[\mathbf{I}(\boldsymbol{\theta})]_{i,j}=-\mathbb{E}_{(T,\mathbf{X}_{T})}\!\left[\frac{\partial^{2}\log f}{\partial\theta_{i}\partial\theta_{j}}\right].$ (17)
We first consider the entry corresponding to the diffusion coefficient $\sigma$. Differentiating (14) with respect to $\sigma$ gives
$\frac{\partial^{2}}{\partial\sigma^{2}}\log f=\frac{D}{\sigma^{2}}-\frac{3}{\sigma^{4}t}\Big{(}(1-t)^{2}+\|\boldsymbol{\Delta}_{t}\|^{2}\Big{)}.$ (18)
From (17), we have
$I_{\sigma\sigma}:=[\mathbf{I}(\boldsymbol{\theta})]_{1,1}$ $=-\mathbb{E}\!\left[\frac{D}{\sigma^{2}}-\frac{3}{\sigma^{4}T}\Big{(}(1-T)^{2}+\|\boldsymbol{\Delta}_{T}\|^{2}\Big{)}\right]$
$=\frac{3}{\sigma^{4}}\,\mathbb{E}\!\left[\frac{(1-T)^{2}}{T}+\frac{\|\boldsymbol{\Delta}_{T}\|^{2}}{T}\right]-\frac{D}{\sigma^{2}}.$ (19)
The two terms in the expectation can be computed separately. Since $T\sim\mathrm{IG}(1,1/\sigma^{2})$ with moments $\mathbb{E}[T]=1$ and $\mathbb{E}[T^{-1}]=1+\sigma^{2}$, it follows that
$\mathbb{E}_{T}\!\left[\frac{(1-T)^{2}}{T}\right]=\sigma^{2}.$ (20)
Moreover, conditioned on $T=t$, the residual vector satisfies $\boldsymbol{\Delta}_{T}\sim\mathcal{N}(0,\sigma^{2}tI_{D-1})$, which gives
$\mathbb{E}_{\mathbf{X}_{T}|T}\!\left[\|\boldsymbol{\Delta}_{T}\|^{2}\,\big{|}\,T\right]=(D-1)\sigma^{2}T.$ (21)
Applying the law of total expectation then yields
$\mathbb{E}\!\left[\frac{\|\boldsymbol{\Delta}_{T}\|^{2}}{T}\right]=\mathbb{E}_{T}\!\left[\frac{1}{T}\,\mathbb{E}_{\mathbf{X}_{T}|T}[\|\boldsymbol{\Delta}_{T}\|^{2}|T]\right]=(D-1)\sigma^{2},$ (22)
and combining the results gives
$I_{\sigma\sigma}=\frac{2D}{\sigma^{2}}.$ (23)
Turning to the cross terms between $\sigma$ and a lateral drift $v^{(k)}$, $k\in\{2,\ldots,D\}$, the mixed derivative is
$\frac{\partial^{2}}{\partial\sigma\partial v^{(k)}}\log f=-\frac{2}{\sigma^{3}}\,\Delta_{t}^{(k)}.$ (24)
From (17), we have
$I_{\sigma,v^{(k)}}:=[\mathbf{I}(\boldsymbol{\theta})]_{1,k}=\frac{2}{\sigma^{3}}\,\mathbb{E}[\Delta_{T}^{(k)}]=0,$ (25)
and similarly, $I_{v^{(k)},\sigma}:=[\mathbf{I}(\boldsymbol{\theta})]_{k,1}=0$, showing that the FIM is diagonal.
Finally, for each drift component $v^{(k)}$, $k\in\{2,\ldots,D\}$, the second derivative takes the form
$\frac{\partial^{2}}{\partial[v^{(k)}]^{2}}\log f=-\frac{t}{\sigma^{2}},$ (26)
which implies
$I_{v^{(k)}v^{(k)}}:=[\mathbf{I}(\boldsymbol{\theta})]_{k,k}=\frac{1}{\sigma^{2}}\,\mathbb{E}[T]=\frac{1}{\sigma^{2}}.$ (27)
In summary, the Fisher information matrix is diagonal, with diffusion contributing $I_{\sigma\sigma}=2D/\sigma^{2}$ and each lateral drift parameter contributing $I_{v^{(k)}v^{(k)}}=1/\sigma^{2}$. This shows that all drift components contribute symmetrically, while the diffusion parameter aggregates information from both temporal and spatial observations.
Figure 1: Empirical conditional joint distribution obtained from particle-based simulations. Here, the drift velocity vector is $(1,0)\leavevmode\nobreak\ \mu\text{m}/\text{s}$.
Figure 2: Analytical conditional joint distribution computed from the closed-form expression. Here, the drift velocity vector is $(1,0)\leavevmode\nobreak\ \mu\text{m}/\text{s}$.
Figure 3: Empirical conditional joint distribution obtained from particle-based simulations under drift velocity vector $(1,-3)\leavevmode\nobreak\ \mu\text{m}/\text{s}$.
Figure 4: Analytical conditional joint distribution derived from the closed-form expression under drift velocity vector $(1,-3)\leavevmode\nobreak\ \mu\text{m}/\text{s}$.
V Numerical Simulations
In this section, we revert to dimensionalized physical parameters to align with common practice in molecular communication studies. All quantities are therefore reported in standard physical units (e.g., micrometers for distance, micrometers per second for velocity, and seconds for time).
We conduct numerical experiments to validate the theoretical channel model in (9). For illustration, we consider the two-dimensional (2D) case with a linear transmitter and a linear absorbing receiver in parallel, and examine the joint distribution of first-hitting time and position.
Particle-based simulations are performed by discretizing Brownian motion with drift via the Euler–Maruyama method [1, Eq. (9)–(11)], and the empirical joint PDF is estimated for comparison. The physical parameters are set as follows: TX–RX distance $\lambda=1\leavevmode\nobreak\ \mu\text{m}$, diffusion coefficient $0.5\leavevmode\nobreak\ \mu\text{m}^{2}/\text{s}$, and initial release position $(0,0)\leavevmode\nobreak\ \mu\text{m}$. Two drift velocity vectors are considered: $(1,0)\leavevmode\nobreak\ \mu\text{m}/\text{s}$ and $(1,-3)\leavevmode\nobreak\ \mu\text{m}/\text{s}$. Each simulation uses $10^{6}$ particles with $2\times 10^{3}$ time steps in the Euler–Maruyama scheme, where each step corresponds to $10^{-3}$ seconds.
Since the simulation horizon is limited to $2$ seconds, the empirical joint PDF should be interpreted as conditional, i.e., conditioned on arrivals within $[0,2]$ seconds. For fair comparison, we likewise compute the corresponding conditional theoretical joint PDF from (9). As shown in Fig. 22 for drift $(1,0)\leavevmode\nobreak\ \mu\text{m}/\text{s}$ (no lateral drift) and Fig. 44 for drift $(1,-3)\leavevmode\nobreak\ \mu\text{m}/\text{s}$ (with lateral drift), the simulation results exhibit excellent agreement with the theoretical predictions.
VI Conclusion
We derived a closed-form expression for the joint distribution of the first arrival time and position of a molecule propagating in a drift–diffusion channel. This result generalizes the classical inverse Gaussian model by explicitly capturing the spatiotemporal coupling of molecular transport in arbitrary spatial dimensions.
Our Fisher information analysis highlights that the joint spatiotemporal observation provides richer information than marginal time-only or position-only models. In particular, lateral drift components, which are not identifiable from first arrival time alone, become estimable when incorporating arrival position. Moreover, the diffusion term aggregates contributions from both temporal and spatial observations, demonstrating the statistical advantage of the joint model for parameter estimation tasks.
These findings suggest that combining temporal and spatial observations could enable more sensitive estimation of key environmental parameters such as background flow direction and diffusivity heterogeneity. Future work may extend this framework to non-uniform drift fields, where drift strength or direction varies spatially.
Acknowledgment
This work was partially supported by the National Science and Technology Council of Taiwan under Grant NSTC 113-2115-M-008-013-MY3. The authors also acknowledge the use of ChatGPT (OpenAI, GPT-5, Aug. 2025) for minor language editing. All technical content and conclusions are the sole responsibility of the authors.